How can you solve an equations with one variable and two numbers if they have a neggative and positive signs.It will be oky if you give me an example of how to solve an equation.If it`s not a problem if you could show step by step please and thank you.
Answer by checkley71(8403)

I'LL TRY.
X+4-3=0 OR X=-4+3 OR X=-1
5X-2+2X-6=0 OR 5X+7X=6+2 OR 12X=8 OR X=8/12 OR X=2/3
7X+6-2X-1=0 OR 7X-2X=-6+1 OR 5X=-5 OR X=-1
DOES THIS GIVE YOU ENOUGH EXAMPLES??